1. The cost of automobile insurance has become a sore subject in California because the rates are dependent on so many variabcontext. 3. Find the approximate p-value for this hypothesis testing. Use inequalities if necessary to answer this question.
Answer #1

Solution:

Given data:

The cost of automobile insurance has become a sore subject in california because the rates are dependents on so many variables

The 6 -month premiums in 2004 for a single male , licensed for 6 to 8 years .

Answer:

We are find out the approximate p- value for the hypothesis testing3. p-value = 2 * P(tdf=3 -1.917) = 0.151 4. 99% CI: (d+5.84sd/n) = (-398.597, 201.597)

PLEASE GIVEME THUMBUP.....